OpenStudy (anonymous):

Find the exact real value of arccosThe quantity square root of two divided by two

OpenStudy (anonymous):

find a number \(x\) where \(\cos(x)=\frac{\sqrt2}{2}\) it should ring a bell

OpenStudy (anonymous):

pi over 4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much (:

OpenStudy (anonymous):

do you know how i can do this Use a calculator to find the approximate value of cos-1(0.55)

OpenStudy (anonymous):

yw make sure when you solve these that you pick the right number in the right range

OpenStudy (anonymous):

depends on the calculator i haven't used one since 2004 just use this http://www.wolframalpha.com/input/?i=arccos%28.55%29 gives answer in both radians and degrees

OpenStudy (anonymous):

I got it ! I just had to type the exact thing on my calculator , and for the first question I asked now its asking me for the exact real value of across The quantity square root of two divided by two , is that one across the actual cos on the unit circle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for arccosine pick the ones on the upper half of the unit circle, i.e. between \(0\) and \(\pi\) for sine pick on the right between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\)

OpenStudy (anonymous):

so is it 11pi over 6 or 7 pi over 4 ? for the cos inverse of root 2 over two ?

OpenStudy (anonymous):

oh no you had it right the first time just \(\frac{\pi}{4}\)

OpenStudy (anonymous):